# The Power of Linear Programming in Sponsored Listings Ranking: Evidence from Field Experiments

### v2

@@ 2.2 Score-based ranking for sponsored listing at Marketplace A
Score-based ranking refers to ranking listed items based on a monotonicity order of a certain score defined by the marketplace. Due to its simplicity and efficiency, score-based ranking is perhaps the most well-used algorithm for recommendation systems and sponsored listing. The key to a score-based ranking is how the score is defined. Score-based ranking is also the production model used in sponsored listing at Marketplace A, and the engineering team has done extensive studies to optimize how the score is defined for improving performance. In this section, we present the high-level ideas of this score-based ranking algorithm used at Marketplace A, which serves as the baseline algorithm in our field study compared with the LP-based algorithm we proposed.
The score-based ranking algorithm deployed for sponsored listing at Marketplace A consists of three major steps: (1) candidate selection, (2) PTR estimation, and (3) ranking, which are detailed below.
@@
Among all the items listed by the sellers, there is only a small number of the items that are related to the seed items and thus are potential candidates for the SLR. The first step is to identify candidate set $\mathcal{N}$. Here, Marketplace A takes two separate approaches. One is based on the categories of seed item, and the other is developed from consumers’ history. Each listed item has been pre-labeled with various categories and sub-categories, spanning multiple layers. The first group of candidate items is determined by their categorical similarity to the seed item. For the second group, Marketplace A trains the k-nearest neighbors (KNN) algorithm featured on each consumer’s browsing and purchase history to select the related items. Eventually, the union of the first and the second group will formulate the candidate set $\mathcal{N}$. Meanwhile, to avoid the overwhelming number of candidate items, Marketplace A will restrict the cardinality of $\mathcal{N}$ during the selection phase. As a result, usually, the number of candidate items $n$ ranges from $1$ to $500$. On the other hand, as indicated previously, the number of slots $m$ might vary among different marketplaces. In general, we consider $m$ to range from $1$ to $50$.
Step 2: purchase-through rate (PTR) estimation.
As Marketplace A adopts the CPA model, the purchase-through rate (PTR) will be regarded as the conversion rate. For each item $j\in\mathcal{N}$, the marketplace would like to predict its PTR, equivalent to the probability of purchase. Naturally, we regard PTR as an appropriate metric to evaluate the relevance of an item, as it reflects the degree to which an item aligns with a consumer’s interest. Using the features of listings (like category, price, and sale history) and consumer preferences (like purchase and browsing history) from historical data, Marketplace A employs a fine-tuned XGBoost model to estimate the PTR. For each item $j\in\mathcal{N}$, there is an associated predicted PTR score, denoted by $r_{j}\in\mathbb{R}$. Further, for each slot $i$, Marketplace A estimates a corresponding weight $h_{i}\in\mathbb{R}$, representing the corresponding slot’s importance. Generally speaking, the first few slots have higher weights than the latter slots.
@@
Step 3: score-based ranking. As the last step, Marketplace A introduces a score, ranks the items based on this score, and then selects the items with the highest scores to display. More specifically, the score $s_{j}$ for the $j$-th item is defined as a weighted sum of the expected organic revenue (PTR $\times$ take rate $\times$ price) and the expected Ad revenue (PTR $\times$ Ad Rate $\times$ price) with the following form:
@@
where $p_{j}$ denotes the price of the item. The term $t_{j}$ denotes the take rate of the item, which is the commission levied by Marketplace A upon the sale of the item, with a predetermined rate for each item category within the marketplace. The organic revenue constitutes the obligatory financial remittance by every vendor for each item sold on the platform. The term $ad_{j}$ denotes the Ad Rate for the $j$-th item, namely, the additional fee vendors are prepared to expend for enhanced item visibility. The Ad revenue represents a discretionary expenditure by sellers aimed at augmenting the prominence of their offerings. The parameter $w\in\mathbb{R}_{+}$ is the weight parameter that controls the balance between organic revenue and ad revenue. Notably, the selection of the scoring function and the calibration of $w$ has gone through extreme tuning at Marketplace over an extended period, and we believe their production model optimally harnesses the power of the score-based algorithm.
3 Linear programming (LP)-based ranking
In this section, we propose an alternative ranking approach based on linear programming (LP) to solve the SLR problem. Specifically, we first present a mixed integer programming (MIP) formulation of the SLR problem. Unfortunately, this MIP problem cannot be solved efficiently within the latency requirement of the online marketplace platform. We then propose a primal-dual algorithm for solving its LP relaxation, which leads to a feasible and almost optimal solution to the original MIP problem.
@@ 3.1 Mixed-Integer Programming (MIP) formulation for the SLR problem
Suppose we have $n$ candidate items designated to be assigned to $m$ slots $(m\leq n)$. For each slot $i$, we have the knowledge of its estimated slot weight $h_{i}$, which represents the importance of the $i$-th slot. In general, the customer is more likely to buy the recommended items at the first few slots than the latter ones, and such difference among slots is measured $h_{i}$. Additionally, for each item $j$, we estimate (1) the relevance to the seed item, $r_{j}$, measured as the purchase rate, and (2) expected revenue $v_{j}$ as the sum of expected organic revenue and Ad revenue as we defined previously from equation (1), i.e., $v_{j}:=r_{j}p_{j}(t_{j}+ad_{j})$. Denote the vectors as $\boldsymbol{v}:=(v_{1},\cdots,v_{n}),\boldsymbol{r}:=(r_{1},\cdots,r_{n}),\boldsymbol{h}:=(h_{1},\cdots,h_{m})$.
Without the loss of generality, we assume that position weights $h_{i}$ are sorted in descending order; namely, the item at the slot with a smaller index is more likely to be purchased than the same item at a slot with the latter index:
@@
In the MIP formulation, we aim to maximize the expected total revenue while posing a lower bound on the overall relevance level. The binary decision variable $\boldsymbol{X}\subseteq\{0,1\}^{m\times n}$ represents a ranking plan, namely, an item $j$ is assigned to slot $i$ when $\boldsymbol{X}_{ij}$ takes the value $1$, or 00 if not. As a feasible ranking plan satisfies $\sum_{i=1}^{m}\boldsymbol{X}_{ij}\leq 1$ (i.e., each slot can only have one candidate), $\sum_{j=1}^{m}\boldsymbol{X}_{ij}\leq 1$ (i.e., each candidate can only be shown at most once), and $\boldsymbol{X}_{ij}\in\{0,1\}$ (i.e., the decision variable is binary). For notational convenience, we denote
@@
$a$ is the maximal possible relevance a ranking can achieve. In the MIP formulation, we aim to find a ranking that maximizes the revenue while making sure that the overall relevance is at least $\lambda$ times the maximal overall relevance with a pre-determined parameter $\lambda\in[0,1]$:
@@
Notice that the MIP formulation is always feasible, because the ranking that gives the maximal overall relevance store is a feasible solution. Additionally, with $\lambda$ decreases, the feasible region of $P_{0}$ is monotonically enlarged. As a result, since a smaller $\lambda$ will induce a possibly lower relevance value but higher revenue, the trade-off between revenue and relevance is fulfilled by the parameter $\lambda$.
@@ 3.2 Linear Programming (LP) relaxation and an efficient algorithm
A challenge of the real-time SLR problem is the latency requirement, which needs to be finished within 0.1 second in general. This poses a strict limitation for obtaining a ranking policy, preventing us from exactly solving MIP. Indeed, the traditional Branch-and-Bound/Branch-and-Cut method for solving MIP costs excessively long and unstable time that the marketplace marketplace cannot afford. In such cases, a typical approach is to look at and solve its LP relaxation. Unfortunately, the state-of-the-art commercial LP solvers can still not solve the LP relaxation within the latency requirement (see Section 3.3 for a computational study). Here, we present a linear time dual-based bisection algorithm that finds a feasible and approximately optimal MIP solution.
@@
=$\mathcal{F}=\left\{\boldsymbol{X}:\sum_{i=1}^{m}\boldsymbol{X}_{ij}\leq 1\forall j,\sum_{j=1}^{n}\boldsymbol{X}_{ij}\leq 1\forall i,\boldsymbol{X}_{ij}\in[0,1]\right\}\ .$ (10)
@@
Suppose we constraint (8) is not redundant, then it follows from complementary slackness that $\mu^{*}>0$. The second step of Algorithm 2 is to find an upper and lower bound on $\mu^{*}$.
@@
Compute $\boldsymbol{X}=\text{PrimalResponse}(\boldsymbol{v},m)$

### v3

@@ 2.3 Score-Based Ranking
Score-based ranking refers to ordering listed items according to the monotonic sequence of a certain score defined by the marketplace. Owing to its simplicity and computational efficiency, score-based ranking is one of the most widely adopted ranking algorithms in both recommender systems and sponsored listings operating under the predict-then-optimize paradigm. The core of this approach lies in how the score is defined. At Marketplace A, score-based ranking is the production model used for sponsored listings, and the engineering team has conducted extensive research to refine the score definition for improved performance. In this section, we present the high-level concept of the score-based ranking algorithm implemented at Marketplace A, which serves as the baseline for comparison with our proposed LP-based algorithm in the field study. Along the way, we also outline the predict-then-optimize process employed by Marketplace A and identify the specific step at which our LP-based algorithm intervenes.
When a consumer opens a webpage with a seed item, Marketplace A first selects a collection of qualified candidate items. From now on, we use the set $\mathcal{N}=\{1,2,...,n\}$ to denote this collection. The number of candidate items $n$ ranges from 1 to 500, as some categories in Marketplace A are well-stocked with many similar items, while others in the long tail have far fewer. Meanwhile, to avoid the overwhelming number of candidate items, Marketplace A restricts the cardinality of $\mathcal{N}$ during the selection phase. Then, Marketplace A is tasked to (1) choose a subset of $\mathcal{N}$ according to the slot number $m(m\leq n)$ and then (2) display this subset of items in ranked order. The slot number $m$ typically ranges from 1 to 50. Consistent with the problem setup and the stringent latency requirements, Marketplace A implemented its SLR procedure through three main stages: (1) candidate selection, (2) purchase-through rate estimation, and (3) ranking, as detailed below.
@@
Among all the items listed on Marketplace A, there is only a small number of items that are related to the seed item and thus are potential candidates for the SLR. The first step is to identify candidate set $\mathcal{N}$. Here, Marketplace A takes two separate approaches. One is based on the categories of seed item, and the other is developed from consumers’ history. Each listed item has been pre-labeled with various categories and sub-categories, spanning multiple layers. The first group of candidate items is determined by their categorical similarity to the seed item. For the second group, Marketplace A trains the k-nearest neighbors (KNN) algorithm featured on each consumer’s browsing and purchase history to select the related items. Eventually, the union of the first and the second group will formulate the candidate set. As observed in the candidate selection stage, the sponsored items can be either complements to the seed item (selected based on the user’s browsing history) or substitutes for the seed item (selected from the same product category).
Step 2: purchase-through rate (PTR) and position weights estimation.
Aligned with common practice in recommender systems, Marketplace A separates the estimation of the conversion rate, namely, the purchase-through rate (PTR), denoted as $r_{j}\in\mathbb{R}$, and the position weights, denoted as $h_{i}\in\mathbb{R}$, from the ranking problem due to latency considerations. In other words, Marketplace A follows the predict-then-optimize paradigm. For the same practical reasons, and consistent with many online marketplaces, a key underlying assumption in solving the SLR problem is that the probability that a consumer clicks on item $j$ at position $i$ can be decomposed as $r_{j}h_{i}$, thereby treating the position weights as independent of the assortment of items. In our study, we take all these estimates as given and intervene only at the ranking stage, which we introduce shortly.
Because Marketplace A adopts a cost-per-acquisition (CPA) model, the key performance metric is the PTR $r_{j}$. Specifically, Marketplace A first generates an offline PTR estimate for each item $j\in\mathcal{N}$, which is then refined using a fine-tuned XGBoost model. This adjustment leverages consumers’ browsing and purchase histories, as well as characteristics of the seed item and other candidate items in $\mathcal{N}$, including category, price, and sales history. Importantly, $r_{j}$ may vary across impressions for the same item $j$, depending on the specific consumer, the seed item, and the other items in the candidate set. An impression refers to an instance in which a user visits a webpage displaying a designated seed item along with a sponsored listing. If the user clicks on a product within the sponsored listing and then returns to the original seed item page, the sponsored listing remains unchanged, and all such activities are considered part of the same impression. In our partner marketplace, PTR also serves as a proxy for the relevance of a sponsored item to the seed item, as it captures the organic purchase likelihood conditional on the seed item [^2] This relevance measure is widely used by our partner marketplace and closely resembles the quality score in Google Ads, which assesses the relevance of an ad and its landing page to a user’s search query: https://support.google.com/google-ads/answer/6167118?hl=en . Decomposing the purchase path in a sponsored listing, the final PTR reflects three main components: the expected click-through rate (CTR), the relevance of the item to the user’s latent purchase intent, and the quality of the landing product page. A key assumption in using PTR as a relevance score is that a click on the seed item reflects the user’s latent purchase intention.. The PTR estimation process also accounts for both the relationship with the seed item and potential cannibalization among items within the sponsored listing. This design reflects a deliberate balance between the latency constraints of real-time operations and the underlying economic relationships among items.
Another important estimation component is the position weights $h_{i}$, which capture the slot-specific probability that a purchase originates from position $i$, independent of the item displayed. These weights are empirically estimated through randomized shuffling and A/B testing. In our empirical analysis, we focus on the desktop website channel, where position weights are fixed across impressions. As expected, the top slots exhibit higher weights than lower ones, and the weights drop sharply once the slot moves to the next page.
Step 3: score-based ranking. In the final step, Marketplace A computes a score for each item, ranks the items based on these scores, and selects those with the highest scores for display. Consequently, not all items in the candidate set appear in the sponsored listing. The number of displayed items is not fixed but determined by Marketplace A’s internal rules. As a result, some sponsored listings may contain no sponsored products, while others may include many.
@@
More specifically, the score $s_{j}$ for the $j$-th item is defined as a weighted sum of the expected organic revenue (PTR $\times$ take rate $\times$ price) and the expected ad revenue (PTR $\times$ ad rate $\times$ price) with the following form:
@@
where $p_{j}$ denotes the price of the item. The term $t_{j}$ denotes the take rate of the item, which is the commission levied by Marketplace A upon the sale of the item, with a predetermined rate for each item category within the marketplace. The organic revenue reflects the mandatory platform fee paid by the marketplace for each transaction. The term $ad_{j}\in\mathbb{R}_{>0}$ denotes the ad rate for the $j$-th item, namely, the amount a vendor is willing to pay for increased item visibility. The process for determining $ad_{j}$ is detailed in Section 2.2. The ad revenue, in contrast to organic revenue, is a discretionary vendor expense aimed at boosting product prominence.
The parameter $w\in\mathbb{R}_{+}$ governs the relative importance of organic revenue versus ad revenue. It is optimized at the impression level based on factors including the predicted PTR, position weights, ad rates, and item prices. The method for optimizing $w$ is detailed in Ge et al., (2021). Notably, the selection of the scoring function and the calibration of $w$ has gone through extreme tuning at Marketplace over an extended period, and we believe their production model optimally harnesses the power of the score-based algorithm.
Next, we will propose an alternative ranking approach based on linear programming (LP) to solve the SLR. Note that the first two steps, candidate selection and PTR and position weights estimation, remain unchanged. Our contribution lies in introducing a new method for the final ranking stage.
3 Linear Programming (LP)-Based Ranking Algorithm
In this section, we first formulate the SLR problem as a mixed integer programming (MIP) model. However, solving this MIP exactly is computationally infeasible under the platform’s latency constraints. We therefore propose a primal–dual algorithm to solve its LP relaxation, which produces a feasible and near-optimal solution to the original MIP. We then describe the production algorithm implemented by Marketplace A, a practical variant of our primal–dual algorithm, that enforces the relevance constraints in expectation rather than at each impression, and present its associated theoretical guarantee. Finally, we provide a theoretical explanation for why our LP-based approach can outperform the conventional score-based algorithm.
@@ 3.1 Mixed-Integer Programming (MIP) Formulation for the SLR Problem
Suppose we have $n$ candidate items designated to be assigned to $m$ slots $(m\leq n)$. For each slot $i$, we have the knowledge of its estimated position weight $h_{i}$, which represents the importance of the $i$-th slot. In general, the customer is more likely to buy the items at the first few slots than the latter ones, and such differences among slots are measured by $h_{i}$. Additionally, for each item $j$, we take the estimates: (1) the relevance to the seed item, $r_{j}$, measured by the PTR, and (2) expected revenue $v_{j}$ as the sum of expected organic revenue and ad revenue as we defined previously from equation (1), i.e., $v_{j}:=r_{j}p_{j}(t_{j}+ad_{j})$. Denote the vectors as $\boldsymbol{v}:=(v_{1},\cdots,v_{n}),\boldsymbol{r}:=(r_{1},\cdots,r_{n}),\boldsymbol{h}:=(h_{1},\cdots,h_{m})$.
Without loss of generality, we assume that position weights $h_{i}$ are sorted in descending order; namely, the item at the slot with a smaller index is more likely to be purchased than the same item at a slot with a larger index.
@@
In the MIP formulation, we aim to maximize the expected total revenue while posing a lower bound on the overall relevance level. The binary decision variable $\boldsymbol{X}\subseteq\{0,1\}^{m\times n}$ represents a ranking plan, namely, an item $j$ is assigned to slot $i$ when $\boldsymbol{X}_{ij}$ takes the value $1$, or 0 if not. As a feasible ranking plan satisfies $\sum_{i=1}^{m}\boldsymbol{X}_{ij}\leq 1$ (i.e., each slot can only have one candidate), $\sum_{j=1}^{n}\boldsymbol{X}_{ij}\leq 1$ (i.e., each candidate can only be shown at most once), and $\boldsymbol{X}_{ij}\in\{0,1\}$ (i.e., the decision variable is binary). For notational convenience, we denote
@@
$a$ is the maximal possible relevance a ranking can achieve. That is, when the ranking is based solely on the relevance (PTR) of an item to the seed item, without accounting for revenue from sponsored products, the maximum achievable relevance score corresponds to the highest possible ranking outcome. In the MIP formulation, we aim to find a ranking that maximizes the revenue while making sure that the overall relevance is at least $\lambda$ times the maximal overall relevance with a pre-determined parameter $\lambda\in[0,1]$:
@@
Notice that the MIP formulation is always feasible, because the ranking that gives the maximal overall relevance store is a feasible solution. Additionally, with $\lambda$ decreases, the feasible region of $P_{0}$ is monotonically enlarged. Since a smaller $\lambda$ will induce a possibly lower relevance value but higher revenue, the trade-off between revenue and relevance is fulfilled by the parameter $\lambda$.
We highlight several key assumptions in our model setup. First, we assume that the position weight $h_{i}$ is independent of the PTR. While this independence may not hold universally across all sponsored listings, the assumption is motivated by the need for low-latency response times. This simplification is commonly adopted in industry practice (Chen et al.,, 2023). Second, while our ranking stage does not explicitly account for interactions among items in a sponsored listing due to latency considerations, we address potential dependencies through PTR estimation. That is, the estimated PTR for each item can depend on the overall assortment of items in the sponsored listing. However, the position weight $h_{i}$ remains fixed and is not influenced by the assortment. This setup reflects a deliberate design choice: we aim to capture item interactions during prediction stage rather than the optimization stage. Third, our model allows for the possibility that a consumer purchases multiple items from a single sponsored listing. This flexibility is consistent with platform objectives, as multiple purchases do not negatively impact the platform, sellers, or consumers. In practice, our partner marketplace data show that, among sponsored listings that led to purchases, most involved only one item, and the maximum number of items purchased from a single listing was four.
@@ 3.2 Linear Programming (LP) Relaxation and an Efficient Algorithm
A challenge of the real-time SLR problem is the latency requirement, which needs to be finished within 0.1 second in general. This poses a strict limitation for obtaining a ranking policy, preventing us from exactly solving MIP. Indeed, the traditional Branch-and-Bound/Branch-and-Cut method for solving MIP costs excessively long and unstable time that the marketplace cannot afford. In such cases, a typical approach is to look at and solve its LP relaxation. Unfortunately, the state-of-the-art commercial LP solvers can still not solve the LP relaxation within the latency requirement (see Appendix A.1 for a computational study). Here, we present a linear time dual-based bisection algorithm that finds a feasible and approximately optimal MIP solution.
@@
$\mathcal{F}=\left\{\boldsymbol{X}:\sum_{i=1}^{m}\boldsymbol{X}_{ij}\leq 1\forall j,\sum_{j=1}^{n}\boldsymbol{X}_{ij}\leq 1\forall i,\boldsymbol{X}_{ij}\in[0,1]\right\}\ .$ (10)
@@
Suppose that constraint (8) is not redundant, then it follows from complementary slackness that $\mu^{*}>0$. The second step of Algorithm 2 is to find an upper and lower bound on $\mu^{*}$.
@@
Compute $\boldsymbol{X}=\text{PrimalResponse}(\boldsymbol{v},m)$ .